Information
Comments
Waited on weather until early morning, while ship slowly made way, checking up on the three drifting moorings that were deployed. Weather settled in time to deploy a CTD cast to 1000m for Si just before dawn, at the same time, the minicosm was rinsed. Steamed to the P-trap location and recovered them, giving the minicosm a 2nd rinse. The O-trap was recovered as the shiny clean minicosm was filled for the next series of experiments. We are now recovering the third and last drifting trap, the I-string. We plan to do a final NIWA cast, and a CTD cast to 800 m for Thorium measurements before departing for station 7 (hopefully) later tonight. Optodes, electrodes, diodes and anodes are being adjusted, replaced and calibrated. Regards Vicky & Karin
Map
A map showing Australia and Antarctica. The map shows the journey of one voyage that has occured in the season, with each route highlighted in a distinct colour.